-five halvs(w-1)=15 what is w??

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

-five halvs(w-1)=15
---------
(-5/2)(w-1)=15
Multiply both sides by -2/5 to get:
w-1=--6
w=-5
Cheers,
Stan H.

Answer by checkley71(8403)

-5/2(W-1)=15
-2.5(W-1)=15
-2.5W+2.5=15
-2.5W=15-2.5
-2.5W=12.5
W=12.5/-2.5
W=-5
PROOF
-2.5(-5-1)=15
-2.5(-6)=15
15=15